OpenStudy (anonymous):

Find all polar coordinates of point P where P = (9, pi/5)

terenzreignz (terenzreignz):

Are you absolutely positive that the question says 'find the POLAR coordinates' ?

terenzreignz (terenzreignz):

oh.... apparently all... :/ okay then

OpenStudy (anonymous):

(9, pi/5 + 2npi) or (-9, pi/5 + 2npi) <- what i think it is Yes it says that :l

terenzreignz (terenzreignz):

Okay, I am only understanding the question now :3 Sorry about that~

OpenStudy (anonymous):

okay sure it's all good

terenzreignz (terenzreignz):

And no :) That's not correct. All the polar coordinates of \(\large (r,\theta)\) would be \[\Large (r\ , \ \theta + 2n\pi) \]and \[\Large (r \ , \ \theta + (2n+1)\pi)\] where \[\Large n \in \mathbb{Z}\]

OpenStudy (anonymous):

OKAY I SEE!!!!! then it will be:

OpenStudy (anonymous):

(9, pi/5 + 2npi) or (-9, pi/5 + (2n+1)pi)

OpenStudy (anonymous):

I got it?

terenzreignz (terenzreignz):

Yeah, that's the one :)

OpenStudy (anonymous):

Yipee! I've got another but ill make a new Q

OpenStudy (anonymous):

Thanks. It's on two pairs of polar coordinates this time